OpenStudy (anonymous):

determine the value of the expressions 1.e. 30% of 60 1.h. 25% of 960 1.k. 80% of 225

OpenStudy (jdoe0001):

\(\bf {\color{brown}{ a}}\%\ of\ b \implies \cfrac{{\color{brown}{ a}}}{100}\cdot b\implies {\color{brown}{ a}}\div 100\cdot b\qquad thus \\ \quad \\ \quad \\ {\color{brown}{ 30}}\%\ of\ 60 \implies \cfrac{{\color{brown}{ 30}}}{100}\cdot 60\implies {\color{brown}{ 30}}\div 100\cdot 60 \\ \quad \\ {\color{brown}{ 25}}\%\ of\ 960\implies \cfrac{{\color{brown}{ 25}}}{100}\cdot 960\implies {\color{brown}{ 25}}\div 100\cdot 960 \\ \quad \\ {\color{brown}{ 80}}\%\ of\ 225 \implies \cfrac{{\color{brown}{ 80}}}{100}\cdot 225\implies {\color{brown}{ 80}}\div 100\cdot 225\)